### v1

@@ 5.2 Limitations and Future Directions
While this study advances the measurement of GenAI literacy in higher education, several limitations warrant attention. Firstly, the GLAT was developed and validated primarily with higher education students, excluding younger K–12 students and educators. Future research should extend its applicability across different educational levels and include educators to ensure broader relevance and utility. Additionally, the study’s examination of incremental validity is based on context-specific tasks involving visual analytics and chatbot interactions. Future investigations should incorporate various contexts and task complexities to comprehensively understand how GenAI literacy affects performance across different domains. The rapidly evolving nature of GenAI technology presents another limitation. As these tools advance, so too must the instruments assessing GenAI literacy. Researchers should continuously update and refine the GLAT to keep pace with new developments and maintain its effectiveness and accuracy (American Educational Research Association et al., 2014). Lastly, it is important to note that the test is conducted in English, which may limit its accessibility and relevance for non-English-speaking participants. Future studies should explore the adaptation of the assessment for different languages to ensure its validity and applicability across diverse linguistic populations.
@@ 6 Conclusion
This study introduced the GLAT, a performance-based instrument designed to assess GenAI literacy within higher education contexts. The GLAT demonstrated robust structural validity and reliability, particularly in evaluating foundational GenAI knowledge among students with varying levels of expertise. The predictive validity of the GLAT further underscored its practical utility, showcasing a significant positive relationship between GLAT scores and learners’ performance in GenAI-supported tasks. This study advocates for the integration of performance-based assessments in addition to traditional self-reported measures to objectively evaluate GenAI literacy. The findings highlight the need for continuous adaptation of assessment tools to keep pace with technological advancements, thereby equipping educators and students with the skills necessary to effectively engage in an AI-driven future. Future research should focus on expanding the applicability of the GLAT across diverse educational levels and contexts, addressing the complex and evolving landscape of GenAI technologies.
Acknowledgement
This study was in part supported by grants from the Australian Research Council (grant agreement numbers DP220101209 and DP240100069 to D.G.). L.Y.’s work is fully funded by the Digital Health CRC (Cooperative Research Centre). D.G.’s work was, in part, supported by the DHCRC and Defense Advanced Research Projects Agency (DARPA) through the Knowledge Management at Speed and Scale (KMASS) program (HR0011-22-2-0047). The DHCRC is established and supported under the Australian Government’s Cooperative Research Centres Program. The U.S. Government is authorised to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of DARPA or the U.S. Government.

### v2

@@ 5.2 Limitations and Future Directions
While this study advances the measurement of GenAI literacy in higher education, several limitations warrant attention. Firstly, the GLAT was developed and validated primarily with higher education students, excluding younger K-12 students and educators. Future research should extend its applicability across different educational levels and include educators to ensure broader relevance and utility. Additionally, the study’s examination of external validity is based on context-specific tasks involving visual analytics and chatbot interactions. Future investigations should incorporate various contexts and task complexities to comprehensively understand how GenAI literacy affects learning performance across different domains. A notable limitation is the focus on certain types of domain knowledge, such as visualisation literacy, without considering other relevant knowledge areas that may affect task performance. Thus, future studies should examine a wider range of domain knowledge to better control its influence on GenAI literacy outcomes. The rapidly evolving nature of GenAI technology presents another limitation. As these tools advance, so too must the instruments assessing GenAI literacy. Researchers should continuously update and refine the GLAT to keep pace with new developments and maintain its effectiveness and accuracy (American Educational Research Association et al., 2014). Lastly, it is important to note that the test is conducted in English, which may limit its accessibility and relevance for non-English-speaking participants. Future studies should explore the adaptation of the assessment for different languages to ensure its validity and applicability across diverse linguistic populations.
@@ 6 Conclusion
This study introduced the GLAT, a performance-based instrument designed to assess GenAI literacy within higher education contexts. The GLAT demonstrated robust structural validity and reliability, particularly in evaluating foundational GenAI knowledge among students with varying levels of expertise. The external validity of the GLAT further underscored its practical utility, showcasing a significant positive relationship between GLAT scores and learners’ performance in GenAI-supported tasks. This study advocates for the integration of performance-based assessments in addition to traditional self-reported measures to evaluate GenAI literacy reliably. The findings highlight the need for continuous adaptation of assessment tools to keep pace with technological advancements, thereby equipping educators and students with the skills necessary to engage in an AI-driven future effectively. Future research should focus on expanding the applicability of the GLAT across diverse educational levels and contexts, addressing the complex and evolving landscape of GenAI technologies.